Prove that the exterior angle of a regular decagon is one - third the interior angle of a regular Pentagon.

Answer by Theo(13342)

the exterior angle of a decagon is equal to 360/10 = 36 degrees.

the internal angle of a pentagon is equal to 180 - 360/5 which is equal to 180 - 72 degrees which is equal to 108 degrees.

3 * 36 is equal to 108 degrees.

the exterior angle of a decagon is equal to 1/3 the interior angle of a pentagon.

these are all regular polygons, by the way.

the formulas only work with regular polygons.

the exterior angle of a pentagon is equal to 360/n, where n is the number of siges.

the interior angle of a polygon can be found by either one of the following formulas:

interior angle = (n-2) * 180 / n

interior angle = 180 - exterior angle.

for the pentagon, the first formula will become 3 * 180 / 5 = 3 * 36 = 108.

for the pentagon, the second formula will becomes 180 - 360/5 = 180 - 72 = 108.

you can take your pick as to which one is easier to work with.

from my perspective, the exterior angle formula is easier to remember.

exterior angle = 360 / n.

the interior angle of a polygon is always supplementary to the exterior angle of the same polygon.

interior angle = 180 - exterior angle.